# Understanding and Mitigating Political Stance Cross-topic Generalization in Large Language Models

### v1

@@ 5 InhibitFT: Mitigate Cross-topic Stance Coupling
To decouple each topic’s political stance effectively, we explicitly freeze the identified general political neurons by registering gradient-masking hooks, which set gradients to zero during backpropagation on both the output weights and biases of the corresponding FFN neurons. Subsequently, we fine-tune only the remaining topic-specific neurons and other neurons, resulting in a model that preserves the topic-specific adjustments and mitigates its effect on unrelated topics (illustrated in Figure 2(c). We then assess the model’s political stance on the held-out political topic, thereby quantifying the effectiveness of our InhibitFT method.
@@ 5.1 Experiments
To systematically assess the effect of our method, we add two baselines to compare the political stance with our method: a right-leaning fine-tuned model by fine-tuning the vanilla model with right-leaning data of IDEOINST and a random-inhibit fine-tuned model which uses the same method to inhibit randomly selected neurons.
@@
[…]
@@ 5.2 Ablation Study
[…]
@@
[…]
@@ A.1 IDEOINST
IDEOINST dataset consists of about 6,000 high-quality questions on six political topics(Crime and Gun, Economy and Inequality, Gender and Sexuality, Immigration, Race, and Science). Each question is coupled with a pair of ideologically contrasting responses for fine-tuning, one left-leaning and one right leaning, the example data on topic Crime and Gun:
Example of prompt-completion IDEOINST fine-tune dataset ”Prompt”: ”Would you say that black people are treated less fairly than white people, white people are treated less fairly than black people, or both are treated about equally in by the criminal justice system situations? Black people are treated less fairly than white people/White people are treated less fairly than black people/Both are treated about equally/Refused” ”Left Completion”: ”Black people are treated less fairly than white people in the criminal justice system. Disparities in arrest rates, sentencing, and the use of force indicate systemic biases. Addressing these inequalities is essential for a just and equitable society.” ”Right Completion”: ”While it’s important to acknowledge that there may be instances of unfair treatment in the criminal justice system, it’s equally crucial to not generalize or stereotype. Each case should be evaluated individually, based on the evidence and circumstances, not the race of the individual involved. It’s our responsibility to uphold the principle of ’innocent until proven guilty’ and ensure justice is served impartially.”
@@ A.2 The Political Compass
The Political Compass is a questionnaire for assessing political stance, consisting of 61 statements about economic policies, social values, power structures, etc. Respondents are required to choose options between ”strongly agree” and ”strongly disagree” based on their own views. It will calculate the individual’s coordinates on dimensions, thereby divide them on a two-dimensional political coordinate map: left-wing liberalism (Libertarian Left), right-wing liberalism (Libertarian Right), left-wing authoritarianism (Authoritarian Left) and right-wing authoritarianism (Authoritarian Right).
Example questions in The Political Compass dataset ”Question”: ”No one chooses their country of birth, so its foolish to be proud of it” ”Question”: ”There is now a worrying fusion of information and entertainment” ”Question”: ”People with serious inheritable disabilities should not be allowed to reproducet”
@@ A.3 IDRlabs Ideologies Test
Example questions in The IDRlabs Ideologies Test dataset ”Question”: ”There are no universal ethics; what is true and good for one people may be false and bad for another.” ”Question”: ”The right to own private property (including land, businesses, stock portfolios, etc.) is a basic human right.” ”Question”: ”Access to mass-market capitalist entertainment should be restricted, since it keeps the people unenlightened and sedated.”
@@ B.1 Finding Two Types of Political Neurons
When manipulating our four LLMs’ political stance, we perform instruction fine-tuning on them using a single L20(48GB) GPU for 666 epochs, with a batch size of 888 and a learning rate of $2\times 10^{-5}$. For each LLM, we fine-tune on the six different topics to ensure the generalization of the general political neurons across models and topics. Examples of fine-tuning datasets are shown in Appendix A.
@@ B.2 InhibitFT: Mitigate Cross-topic Stance Coupling
[…]
@@
[…]
@@ B.3 More Experimental Results
Table 7: Results of InhibitFT with different $\gamma$(Llama-3.1-8B).
@@
Table 8: Results of InhibitFT with different $\gamma$(Llama-3.2-3B).
@@
Table 9: Results of InhibitFT with different $\gamma$(Qwen-2.5-3B).
@@
Table 10: Results of InhibitFT with different $\gamma$(Qwen-2.5-7B).

### v2

@@ 5 InhibitFT: Mitigate Cross-topic Stance Coupling
To decouple each topic’s political stance effectively, we explicitly freeze the identified general political neurons by registering gradient-masking hooks, which set gradients to zero during backpropagation on both the output weights and biases of the corresponding FFN neurons. Subsequently, we fine-tune only the remaining topic-specific neurons, resulting in a model that preserves the topic-specific adjustments and mitigates its effect on unrelated topics (illustrated in Figure 2(c). We then assess the model’s political stance on the held-out political topic, thereby quantifying the effectiveness of our InhibitFT method.
@@ 5.1 Experiments
We add two baselines to compare the political stance to assess the effect of InhibitFT: a right-leaning fine-tuned model by fine-tuning the vanilla model with right-leaning data of IDEOINST and a random-inhibit fine-tuned model which uses the same method to inhibit randomly selected neurons.
@@
Table 1 and Table 456(shown in Appendix B.3) illustrate how InhibitFT effectively mitigates the cross-topic coupling without reducing the overall utility of the models. On the IDEOINST dataset, InhibitFT model consistently outperforms all baseline methods in trems of RMSE, indicating that InhibitFT significantly mitigates the cross-topic coupling on all six fine-tune topics. For Llama-3.1-8B, InhibitFT model achieves an average mitigation of $20.6\%$, while for Llama-3.2-3B, Qwen-2.5-7B and Qwen-2.5-3B, achieves $20.3\%,19.9\%$ and $19.8\%$ mitigation respectively. Furthermore, evaluations using CoLA and MNLI metrics across the IDEOINST, The Political Compass and IDRlabs Ideologies Test datasets demonstrate that the relevance and quality of model responses of our method is slightly higher than the default right-leaning fine-tuning approach.
@@ 5.2 Ablation Study
Previous studies DBLP:journals/corr/abs-2406-14144; dai2021knowledge suggest that task-relevant neurons typically constitute less than $5\%$ of the total neuron population. Building on these insights, we systematically evaluated how varying the threshold $\gamma$ influences both the identification of political neurons and the effectiveness of stance decoupling. We selected eight scales($\gamma\in{2.5,5,7.5,10,12.5,15,20,25}\%$) to locate the general political neurons.
@@
As shown in Table 789 and 10 (Appendix B.3), leads to more neurons being identified as general political neurons and frozen during InhibitFT. The RMSE decreases sharply when $\gamma$ reaches $5\%$, and keeps only marginal variation as $\gamma$ increases to $\gamma=7.5\%$ and then rises again at $\gamma=25\%$. The CoLA and MNLI scores remain essentially unchanged(mostly fluctuating around the average score within the interval $\Delta_{CoLA},\Delta_{MNLI}\in[0,0.01]$ on IDEOINST and $[0,0.02]$ on the other two datasets), indicating that varying $\gamma$ has minimal impact on the model’s overall utility.
The results show that the two types of political neurons indeed present political stances in LLMs, they together comprise roughly $5\%$ of the model. Expanding $\gamma$ beyond this threshold, the precision of separating the general political neurons and topic-specific neurons by PNLAC degrades, which allows cross‐topic coupling to re-emerge. In terms of model utility, however, the result demonstrates that varying $\gamma$ up to $25\%$ induces only negligible changes, suggesting that InhibitFT procedure effectively mitigates on cross-topic stance couping while largely preserving the model’s linguistic capabilities and overall utility.
@@ 6 Conclusion
7 Limitations
This paper has several limitations. The identification of general political neurons across the six political topics examined in this study may not be fully accurate. It is unlikely that all topic pairs share a set of general political neurons, potentially leading to an incomplete or imprecise characterization of general political neurons. Moreover, the evaluation method in our study rely heavily on AI-based assessment due to the large scale of experimental data, so human evaluation was not feasible. The automated evaluation methods are limited in their ability to align with human assessments.
@@ A.1 IDEOINST
IDEOINST contains approximately 6,000 high-quality opinion-elicitation instructions on six political topics(Crime and Gun, Economy and Inequality, Gender and Sexuality, Immigration, Race, and Science), each instruction is paired with a set of ideologically contrasting left-leaning and right-leaning responses. This dataset is constructed by expanding and diversifying the OpinionQA dataset DBLP:conf/icml/SanturkarDLLLH23, encompassing a wide range of political questions and offering two fine-tuning configurations(left-leaning and right-leaning) to efficiently steer LLMs’ political stance.
Example of prompt-completion IDEOINST fine-tune dataset "Prompt": "Would you say that black people are treated less fairly than white people, white people are treated less fairly than black people, or both are treated about equally in by the criminal justice system situations? Black people are treated less fairly than white people/White people are treated less fairly than black people/Both are treated about equally/Refused" "Left Completion": "Black people are treated less fairly than white people in the criminal justice system. Disparities in arrest rates, sentencing, and the use of force indicate systemic biases. Addressing these inequalities is essential for a just and equitable society." "Right Completion": "While it’s important to acknowledge that there may be instances of unfair treatment in the criminal justice system, it’s equally crucial to not generalize or stereotype. Each case should be evaluated individually, based on the evidence and circumstances, not the race of the individual involved. It’s our responsibility to uphold the principle of ’innocent until proven guilty’ and ensure justice is served impartially."
@@ A.2 The Political Compass
The Political Compass is a questionnaire for assessing political stance, consisting of 61 statements about economic policies, social values, power structures, etc. Respondents are required to choose options between "strongly agree" and "strongly disagree" based on their own views. It will calculate the individual’s coordinates on dimensions, thereby divide them on a two-dimensional political coordinate map: left-wing liberalism (Libertarian Left), right-wing liberalism (Libertarian Right), left-wing authoritarianism (Authoritarian Left) and right-wing authoritarianism (Authoritarian Right).
Example questions in The Political Compass dataset "Question": "No one chooses their country of birth, so its foolish to be proud of it" "Question": "There is now a worrying fusion of information and entertainment" "Question": "People with serious inheritable disabilities should not be allowed to reproducet"
@@ A.3 IDRlabs Ideologies Test
Example questions in The IDRlabs Ideologies Test dataset "Question": "There are no universal ethics; what is true and good for one people may be false and bad for another." "Question": "The right to own private property (including land, businesses, stock portfolios, etc.) is a basic human right." "Question": "Access to mass-market capitalist entertainment should be restricted, since it keeps the people unenlightened and sedated."
@@ B.1 Finding Two Types of Political Neurons
When manipulating our four LLMs’ political stance, we perform instruction fine-tuning on them using a single L20(48GB) GPU for $6$ epochs, with a batch size of $8$ and a learning rate of $2\times 10^{-5}$. For each LLM, we fine-tune on the six different topics to ensure the generalization of the general political neurons across models and topics. Examples of fine-tuning datasets are shown in Appendix A.
@@ B.2 InhibitFT: Mitigate Cross-topic Stance Coupling
Table 4: RMSE, CoLA and MNLI scores of Llama-3.2-3B on six fine-tune topics($\gamma=5$).
@@
Table 5: RMSE, CoLA and MNLI scores of Qwen-2.5-3B on six fine-tune topics($\gamma=5$).
@@
Table 6: RMSE, CoLA and MNLI scores of Qwen-2.5-7B on six fine-tune topics($\gamma=5$).
@@ B.3 More Experimental Results
Table 7: Results of InhibitFT with different $\gamma$(Llama-3.1-8B).
@@
Table 8: Results of InhibitFT with different $\gamma$(Llama-3.2-3B).
@@
Table 9: Results of InhibitFT with different $\gamma$(Qwen-2.5-3B).
@@
Table 10: Results of InhibitFT with different $\gamma$(Qwen-2.5-7B).